Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
6x + 2y = 2
3x + 5y = 5

Answer by checkley77(12844)

6x + 2y = 2
3x + 5y = 5 multiply this equation by -2 & then add them.
-6x-10y=-10
6x+2y=2
-----------------
-8y=-8
y=-8/-8
y=1 ans.
6x+2*1=2
6x+2=2
6x=2-2
6x=0
x=0 ans.
Proof:
3*0+5*1=5
5=5